# Dynamic linear regression models for forecasting time series with semi long memory errors

### v1

@@ 3.3. Parameterisations and prior distributions
The autoregressive parameters $\phi_{i}$ are reparameterised in terms of partial autocorrelations $-1\leq\widetilde{\phi}_{i}\leq 1$, for $i=1,\dots,p$, following Barndorff-Nielsen and Schou, (1973) to ensure the process is always stationary. For the moving average parameters $\psi_{j}$, for $j=1,\dots,q$, the same reparameterisation in terms of $\widetilde{\psi}_{j}$ ensures the process is always invertible. To make the variance parameter $\sigma^{2}_{\varepsilon}$ and the tempering parameter $\lambda$ unrestricted, we reparameterise using log-transformations. For the ARTFIMA process, the fractional parameter $d$ ($\notin\mathbb{Z}$) is unrestricted (since stationarity does not depend on $d$). For ARFIMA models, on the other hand, $-0.5<d<0.5$ is a necessary condition for stationarity. We thus reparameterise $d$ to an unrestricted parameter $\widetilde{d}$ by a scaled Fisher transformation, i.e. $\widetilde{d}=\mathrm{arctanh}(2d)$.
@@
When proposing a sample from the multivariate normal proposal as described in Section 3.2, a prior on the interval $[-1,1]$ for each autoregressive and moving average parameter ensures $|\widetilde{\phi}_{i}|\leq 1$ and $|\widetilde{\psi}_{j}|\leq 1$ for the $i$ and $j$ above. We set independent (within and between) uniform priors $\boldsymbol{\phi}\sim\mathrm{Unif}(-1,1)^{p}$ and $\boldsymbol{\psi}\sim\mathrm{Unif}(-1,1)^{q}$. The priors for $\log(\sigma^{2}_{\varepsilon})$ and $\log(\lambda)$ are $\mathcal{N}(0,100)$. For the ARTFIMA process $d\sim\mathcal{N}(0,1)$, while for ARFIMA models $\widetilde{d}\sim\mathcal{N}(0,1)$ which implies a weakly informative on $(-0.5,0.5)$ for $d$. Finally, $\beta_{k}$, for $k=1,\dots,m$, are assigned independent $\mathcal{N}(0,100)$ priors. We have verified that our results are not influenced by the choice of prior (the datasets studied are very large).
4. Simulated examples
4.1. Approximating the time domain likelihood
@@
We compare posterior inference via the Whittle likelihood to those of using the exact Gaussian and Kalman filter likelihoods under the same weakly informative priors for dynamic linear regression models on simulated data. By the exact Gaussian likelihood, we mean evaluating a multivariate normal density where the covariance matrix is computed from the autocovariance function. The Kalman filter likelihood, on the other hand, performs the Kalman filter recursions to compute the likelihood. Unlike the Gaussian likelihood approach, the Kalman filter avoids inverting a large covariance matrix. Note that both these methods use the same (exact) time domain likelihood, however, evaluated in a different manner. This example is restricted to an ARMA process as the ARFIMA/ARTFIMA likelihoods cannot be efficiently evaluated using the Kalman filter.
@@
Let $y_{t}=\beta_{1}x_{t}+\eta_{t}$ where $\eta_{t}$ is an ARMA(3,1) model with $T=10{,}001$. We set $\beta_{1}=3$ and the reparameterised vector
$\boldsymbol{\vartheta}=\left(\widetilde{\phi}_{1},\widetilde{\phi}_{2},\widetilde{\phi}_{3},\widetilde{\psi}_{1},\log(\sigma_{\varepsilon}^{2})\right)^{\top}=\left(0.4,-0.2,0.1,-0.2,\log(2)\right)^{\top}.$
@@
The exogenous predictor $x_{t}$ is assumed to be known and is simulated from an ARMA(2,2) process. MCMC is used to sample the three posteriors via $10{,}000$ iterations of the random walk Metropolis-Hastings algorithm. Figure 2 plots the kernel density estimates of the marginal posteriors for each (untransformed) parameter and shows that the three posteriors are almost identical. Hence, the $5{,}000$ ($(T-1)/2$) periodogram ordinates are sufficient to give a very accurate approximation of the true time domain likelihood in this case.
@@
Figure 2. Kernel density estimates of the marginal Whittle, Kalman filter and Gaussian posteriors for a DLR model with $\eta_{t}\sim$ ARMA(3,1).
@@
Figure 3 reports the effective sample sizes. The figure shows that the Gaussian likelihood has a $\approx 25\%$ higher ESS for the $\beta_{1}$ parameter compared to the Whittle likelihood and the Kalman filter. The Whittle approach has a lower ESS for $\sigma^{2}$ compared to its time domain counterparts. The AR and MA parameters are comparable for all three posteriors, with the Whittle posterior having slightly higher ESS for all of them.
@@
Figure 3. Effective sample sizes for $10{,}000$ MCMC iterations for each parameter in a DLR model with $\eta_{t}\sim\mathrm{ARMA}(3,1)$. The results are shown for the three methods in the legend.
@@
Table 1 presents the run times (in seconds) of each method. Our proposed Whittle approach is roughly 8 times faster than the Gaussian likelihood and 94 times faster than the Kalman filter. The Gaussian likelihood uses the SuperGauss package in R (Ling and Lysy,, 2017), which scales as $\mathcal{O}(T\log_{2}^{2}(T))$ from the superfast Toeplitz algorithm proposed in Ammar and Gragg, (1988). Despite the Kalman filtering being performed in $\mathcal{O}(T)$ operations, the algorithm sweeps through all observations recursively for each new MCMC proposal and is thus time-consuming, especially in the Python implementation used in this paper. The Whittle log-likelihood (3.7), on the other hand, is a simple sum that only requires computing $I_{Z}(\omega)$ in (3.3) and the relevant spectral density, and is thus much faster than the Kalman filter.
@@
Table 1. Computation time for sampling $10{,}000$ MCMC draws for each method.
Figure 4. Quantile-quantile (QQ) plots of the ratio $I_{Z}(\omega_{k})/f(\omega_{k})$ for simulated ARTFIMA vs ARFIMA models for the three lowest positive frequencies. The top row, middle row, and bottom row, correspond to $T=1001$, $T=10001$, and $T=20001$, respectively.
@@ 4.2. The periodogram distribution for low frequencies
This section empirically verifies the periodogram distribution for small frequencies for the ARTFIMA process. We also illustrate how this fails for the ARFIMA process. The latter has been investigated theoretically in Robinson, (1995); Rousseau et al., (2012) and empirically in Meerschaert et al., (2014); Sabzikar et al., (2019).
We test both models without the regression coefficient, i.e. $\boldsymbol{\beta}=\boldsymbol{0}$, and using
  1. (M1)
    ARTFIMA$(2,d,\lambda,0)$ with $(\phi_{1},\phi_{2},d,\lambda,\sigma^{2})=(0.742,0.227,2.139,0.616,1)$,
  2. (M2)
    ARFIMA$(2,d,0)$ with $(\phi_{1},\phi_{2},d,\sigma^{2})=(1.466,-0.525,0.493,1)$.
@@
These parameter values are chosen to match those obtained by the real data application in Section 5.5. It follows from (3.5) that the ratio of the periodogram and its spectrum follows the distribution
(4.1) $\frac{I_{Z}(\omega_{k})}{f(\omega_{k})}\sim\frac{\chi^{2}(2)}{2},$
for $\omega_{k}\in\Omega$ as $T\rightarrow\infty$.
@@
We verify (4.1) by simulating $10{,}000$ realisations for $T=1{,}001,10{,}001,20{,}001$ and keeping the three smallest positive Fourier frequencies $\omega_{k}$ (excluding zero). Figure 4 displays the quantile-quantile (QQ) plots of the simulated quantiles vs theoretical quantiles of (4.1) for all three cases of $T$. The left and right panels display the plots for the ARTFIMA and ARFIMA models, respectively. The upper, middle, and bottom rows correspond to $T=1{,}001,10{,}001,20{,}001$, respectively. The left panel of Figure 4 shows that for the three lowest frequencies, the empirical quantiles of the periodogram ratios are almost identical to their theoretical counterpart for all $T$. This confirms the suitability of (4.1) for the small frequencies in the ARTFIMA model case. In contrast, the right panel shows that (4.1) is not suitable for the small frequencies in the ARFIMA model case, and that an increasing $T$ does not make (4.1) more suitable.
@@
The violation of (4.1) is partly attributed to the fact that the spectrum of the ARFIMA model behaves as a power law and diverges as $\omega\rightarrow 0$. Sabzikar et al., (2019) and Meerschaert et al., (2014) show for applications such as hydrology, finance, and climatology, the ARFIMA spectral density provides a poor fit of the periodogram at the low frequencies. We show the same issue in a dynamic linear regression model with ARFIMA errors; see Figure 10 in Section 5.4. A pragmatic solution is to use a low-frequency cutoff in the estimation of ARFIMA models in the frequency domain, however, this incurs a substantial loss of information in the estimation of the long-memory parameter $d$. Section 5 thus uses the exact Gaussian likelihood for dynamic linear models with ARFIMA errors, which corresponds to the exact maximum likelihood approach in Doornik and Ooms, (2004). The resulting computation is much slower than a frequency domain approach and stresses the importance of our contribution, namely: A computationally fast dynamic linear model that accounts for long-range dependence in the error process.
@@ 5.1. Overview and objectives
This section analyses electricity demand as the response variable conditional on relevant exogenous variables for two real-world datasets. The two datasets contain data for New England, United States of America (Example 1), and Victoria, Australia (Example 2), presented in Sections 5.4 and 5.5, respectively. Forecasting Example 2 is considerably more challenging than Example 1 and we show that accounting for long-range dependence is essential to obtain good forecast accuracy for moderately long horizons. In contrast, long-range dependency does not improve the forecasts in Example 1, however, we show it has a significant impact on the posterior distribution of the regression parameters $\boldsymbol{\beta}$.
@@ 5.2. Model settings
Given an error process (ARTFIMA, ARFIMA, or ARMA) we select the orders of $p$ and $q$ via the smallest Bayesian information criterion (BIC) value (Schwarz,, 1978). We restrict $0\leq p\leq 3$ and $0\leq q\leq 2$ for Example 1, and $0\leq p\leq 2$ and $0\leq q\leq 1$ for Example 2. The reason for smaller lag orders for the second example is that this example includes additional seasonal lags. After a given order $p,q$ for an error process is determined, we compare its forecasting ability to the two remaining error processes using the same $p,q$. We do this for all three error processes, so that each one has the opportunity to forecast using its optimal lag orders (the lag order may be the same for different error processes).
@@
It is well known that the log-likelihood function for ARMA models may exhibit multimodality, and this is also true for ARTFIMA and ARFIMA. To avoid finding a poor local mode, we use global optimisation via Basin-hopping (Li and Scheraga,, 1987) to find the maximum likelihood estimate when computing the BIC.
5.3. Forecasting evaluation settings
For a given $p$ and $q$ determined as described in Section 5.2, the error processes are compared with respect to their forecasting abilities in a time series cross-validation setting. We use $k=100$ out-of-sample observations with different forecast horizons, $h=1,\dots,15$. The exogenous variables $\boldsymbol{X}_{t}$ are not modeled, i.e. their observed values are used for any forecasting horizon. A model to forecast future values of $\boldsymbol{X}_{t}$ is needed in practice, however, this simplifying assumption is made for all cases and does not affect the relative performance.
We evaluate both point and distributional forecasts. The optimal $h$-step ahead point forecast under a squared loss function is the conditional expectation
(5.1) $\mu_{y^{(h)}_{T+i}|\mathcal{I}_{T+i-h}}=\mathrm{E}(y_{T+i}|\mathcal{I}_{T-h-i}),$
where $\mathcal{I}_{b}$ denotes the information set at $t=b$, i.e. all the variables known up to and including time $t=b$. We evaluate the accuracy of the point forecast via the (out-of-sample) root mean squared error (RMSE)
$\text{RMSE}^{(h)}=\sqrt{\sum^{k}_{i=1}\frac{\left(y_{T+i}-\widehat{y}^{(h)}_{T+i}\right)^{2}}{k}},$
where $\widehat{y}^{(h)}_{T+i}$ is an estimate of (5.1) obtained using the maximum a posteriori (MAP) estimate for each model. To assess distributional forecasts, we obtain the $h$-step ahead predictive posterior density

### v2

@@ 3.3. Parameterisations and prior distributions
The autoregressive parameters $\phi_{i}$ are reparameterised in terms of partial autocorrelations $-1<\widetilde{\phi}_{i}<1$, for $i=1,\dots,p$, following Barndorff-Nielsen and Schou, (1973) to ensure the process is always stationary. For the moving average parameters $\psi_{j}$, for $j=1,\dots,q$, the same reparameterisation in terms of $\widetilde{\psi}_{j}$ ensures the process is always invertible and the moving average parametrisation is unique. To make the variance parameter $\sigma^{2}_{\varepsilon}$ and the tempering parameter $\lambda$ unrestricted, we reparameterise using log-transformations. For the ARTFIMA process, the fractional parameter $d$ ($\notin\mathbb{Z}$) is unrestricted (since stationarity does not depend on $d$). For ARFIMA models, on the other hand, $-0.5<d<0.5$ is a necessary condition for stationarity. We thus reparameterise $d$ to an unrestricted parameter $\widetilde{d}$ by the scaled Fisher transformation $\widetilde{d}=\mathrm{arctanh}(2d)$.
@@
When using a sample from the multivariate normal proposal as described in Section 3.2, a prior on the interval $(-1,1)$ for each autoregressive and moving average parameter ensures $|\widetilde{\phi}_{i}|<1$ and $|\widetilde{\psi}_{j}|<1$ for the $i$ and $j$ above. We set independent (within and between) uniform priors $\boldsymbol{\phi}\sim\mathrm{Unif}(-1,1)^{p}$ and $\boldsymbol{\psi}\sim\mathrm{Unif}(-1,1)^{q}$. The priors for $\log(\sigma^{2}_{\varepsilon})$ and $\log(\lambda)$ are $\mathcal{N}(0,100)$. For the ARTFIMA process $d\sim\mathcal{N}(0,1)$, while for ARFIMA models, the prior $\widetilde{d}\sim\mathcal{N}(0,1)$ implies a weakly informative on $(-0.5,0.5)$ for $d$. Finally, $\beta_{k}$, for $k=1,\dots,m$, are assigned independent $\mathcal{N}(0,100)$ priors. We have verified that our results are not influenced by the choice of prior (because the datasets studied are very large).
4. Simulation studies
4.1. Accuracy of the approximation of the time domain likelihood
@@
We compare posterior inference via the Whittle likelihood to those using the exact Gaussian and Kalman filter likelihoods under the same weakly informative priors for dynamic linear regression models on simulated data. By the exact Gaussian likelihood, we mean evaluating a multivariate normal density where the covariance matrix is computed from the autocovariance function. The Kalman filter likelihood performs the analytic filtering steps recursively to compute the likelihood (Hamilton,, 2020). Unlike the Gaussian likelihood approach, the Kalman filter avoids inverting a large covariance matrix. Note that both methods use the same (exact) time-domain likelihood, but are evaluated differently. This example is restricted to an ARMA process as the ARFIMA/ARTFIMA likelihoods cannot be efficiently evaluated using the Kalman filter (recall from Section 1 that they do not have finite-dimensional state space representations).
@@
Let $y_{t}=\beta_{1}x_{t}+\eta_{t}$ where $\eta_{t}$ is an ARMA(3,1) model with $T=5{,}001$. We simulate using the true value $\beta_{1}=\beta^{(0)}_{1}=3$ and the reparameterised vector ($\psi_{1}^{(0)}=\widetilde{\psi}_{1}^{(0)}$)
$\boldsymbol{\vartheta}^{(0)}=\left(\widetilde{\phi}_{1}^{(0)},\widetilde{\phi}_{2}^{(0)},\widetilde{\phi}_{3}^{(0)},\psi_{1}^{(0)},\log(\sigma_{\varepsilon}^{2(0)})\right)^{\top}=\left(0.4,-0.2,0.1,0.2,\log(2)\right)^{\top},$
and with $\left(\phi_{1}^{(0)},\phi_{2}^{(0)},\phi_{3}^{(0)}\right)^{\top}=\left(0.5,-0.248,0.1\right)^{\top}$ in the original parameterisation, where the superscript (0) denotes a true value. The exogenous predictor $x_{t}$ is assumed to be known and is simulated from an ARMA(2,2) process. The MCMC algorithm described in Section 3.2 is used to sample the three posteriors for $10{,}000$ iterations with a burn-in of $3{,}000$ samples. Table 1 reports the mean squred error (MSE) of the posterior mean for $1{,}000$ replicates of the data for each parameter. The MSE of both the Whittle likelihood and the Gaussian likelihood are small and the differences are likely attributed to Monte Carlo error.
MSE $\beta_{1}$ $\phi_{1}$ $\phi_{2}$ $\phi_{3}$ $\theta_{1}$ $\sigma^{2}$
Gaussian 0.013 0.076 0.007 0.003 0.072 0.002
Whittle 0.014 0.065 0.006 0.002 0.063 0.002
Table 1. The mean square error (MSE) for posterior means (reparameterised) with the Gaussian likelihood and Whittle likelihood under the same prior for $1{,}000$ data simulation replicates.
@@
Figure 2 plots kernel density estimates of the marginal posteriors for each parameter given data from one simulation. As shown, the three posteriors are almost identical. Hence, the $n=2{,}500$ periodogram ordinates were enough to get a very close approximation of the Whittle likelihood to the true time domain likelihood.
@@
Figure 2. Kernel density estimates of the marginal Whittle, Kalman filter and Gaussian posteriors (original parameterisation) for a DLR model with $\eta_{t}\sim$ ARMA(3,1).
@@
Figure 3 reports the effective sample sizes. The figure shows that the Gaussian likelihood has a $\approx 25\%$ higher ESS for the $\beta_{1}$ parameter compared to the Whittle likelihood and the Kalman filter likelihood. The Whittle approach has a lower ESS for $\sigma^{2}$ compared to its time domain counterparts. The AR and MA parameters are comparable for all three posteriors, with the Whittle posterior having slightly higher ESS for all.
@@
Figure 3. Effective sample sizes for $10{,}000$ MCMC iterations for each parameter (original parameterisation) in a DLR model with $\eta_{t}\sim\mathrm{ARMA}(3,1)$. The results are shown for the 3 methods in the legend.
@@
Table 2 presents the run times (seconds) of each method. The proposed Whittle approach is roughly 8 times faster than the Gaussian likelihood and 94 times faster than the Kalman filter. The Gaussian likelihood uses the SuperGauss package in R (Ling and Lysy,, 2017), which scales as $\mathcal{O}(T\log_{2}^{2}(T))$ using the superfast Toeplitz algorithm proposed in Ammar and Gragg, (1988). Despite the Kalman filtering being performed in $\mathcal{O}(T)$ operations, the algorithm sweeps through all observations recursively for each new MCMC proposal and is thus time-consuming, especially in the Python implementation used in this paper. The Whittle log-likelihood (3.9), on the other hand, is a simple sum that only requires computing $I_{Z}(\omega)$ in (3.5) and the relevant spectral density, and is thus much faster than the Kalman filter.
@@
Table 2. Computation time (seconds) for sampling $10{,}000$ MCMC draws for each method.
Figure 4. Quantile-quantile (QQ) plots of the ratio ($\widehat{\beta}_{1}$ varying over replicates) $I_{Z}(\omega_{k};\widehat{\beta}_{1})/f(\omega_{k};\boldsymbol{\vartheta}^{(0)})$ for simulated DLR ARTFIMA (left panel) vs DLR ARFIMA (right panel) models for the three lowest positive frequencies. The top row, middle row, and bottom row, correspond to $T=1{,}001$, $T=10{,}001$, and $T=20{,}001$, respectively.
@@ 4.2. The periodogram distribution for low frequencies
This section empirically verifies the periodogram distribution for small frequencies for the DLR with ARTFIMA errors. We also illustrate how this fails in the case of an ARFIMA error process. In a non-DLR setting, i.e. $\boldsymbol{\beta}=\mathbf{0}$, the latter has been investigated theoretically in Robinson, (1995); Rousseau et al., (2012) and empirically in Meerschaert et al., (2014); Sabzikar et al., (2019). The aim of this simulation study is to empirically verify that the ARTFIMA scaled low-frequency periodogram observations of the pseudo data in (3.5) follows (3.6), and that it is violated in the ARFIMA case.
Consider the models $y_{t}=\beta_{1}x_{t}+\eta^{(i)}_{t}$ for $i=1,2$,
$\eta^{(1)}\sim\mathrm{ARTFIMA}(2,d,\lambda,0),\,\,\eta^{(2)}\sim\mathrm{ARFIMA}(2,d,\lambda,0),$
[…]
  1. $(i=1)$
    $\boldsymbol{\vartheta}^{(0)}=(\phi_{1}^{(0)},\phi_{2}^{(0)},d^{(0)},\lambda^{(0)},\sigma^{2(0)})^{\top}=(0.742,0.227,2.139,0.616,1)^{\top}$,
  2. $(i=2)$
    $\boldsymbol{\vartheta}^{(0)}=(\phi_{1}^{(0)},\phi_{2}^{(0)},d^{(0)},\sigma^{2(0)})^{\top}=(1.466,-0.525,0.493,1)^{\top}$.
@@ 2. $(i=2)$
The parameter values of the error processes are chosen to match those obtained by the real data application in Section 5.5. Recall from (3.6) that the ratio of the periodogram and its spectrum follows the distribution
$\frac{I_{Z}(\omega_{k};\beta_{1})}{f(\omega_{k};\boldsymbol{\vartheta})}\sim\frac{\chi^{2}(2)}{2},$ (4.1)
@@
as $T\rightarrow\infty$ for $\omega_{k}=\frac{2\pi k}{T},k\in\mathcal{K}$, where $\mathcal{K}$ is defined in (3.2) and the dependence of the parameters is emphasised. Note that if we set $\beta_{1}=\beta_{1}^{(0)}$ when computing the periodogram, then this corresponds to the known results in a non-DLR setting since the pseudo data in (3.1) by the construction of the simulation follows an ARTFIMA (or ARFIMA) process. To obtain $\beta_{1}\neq\beta_{1}^{(0)}$, we evaluate the periodogram at the MAP $\widehat{\beta}_{1}$ in each replication. We verify (4.1) with $\boldsymbol{\vartheta}=\boldsymbol{\vartheta}^{(0)}$ and $\beta_{1}=\widehat{\beta}_{1}$ by simulating $10{,}000$ realisations for $T=1{,}001,10{,}001,20{,}001$ and keeping the three smallest positive Fourier frequencies $\omega_{k}$ (excluding zero). Figure 4 displays the quantile-quantile (QQ) plots of $I_{Z}(\omega_{k};\widehat{\beta}_{1})/f(\omega_{k};\boldsymbol{\vartheta}^{(0)})$ against $\chi^{2}_{2}/2$ for all three cases of $T$. The left and right panels display the plots for the ARTFIMA and ARFIMA models, respectively. The upper, middle, and bottom rows correspond to $T=1{,}001$, $T=10{,}001$ and $T=20{,}001$, respectively. The left panel (ARTFIMA) of Figure 4 shows that for the three lowest frequencies, the empirical quantiles of the periodogram ratios are almost identical to their theoretical counterpart for all $T$. This confirms the suitability of (4.1) for the small frequencies in the ARTFIMA model case. In contrast, the right panel (ARFIMA) shows that (4.1) is unsuitable for the small frequencies in the ARFIMA model case, and that an increasing $T$ does not make (4.1) more suitable.
@@
The violation of (4.1) is partly attributed to the fact that the spectrum of the ARFIMA model behaves as a power law and diverges as $\omega\rightarrow 0$. Sabzikar et al., (2019) and Meerschaert et al., (2014) show that for applications such as hydrology, finance, and climatology, the ARFIMA spectral density provides a poor fit of the periodogram at the low frequencies. We show the same issue in a dynamic linear regression model with ARFIMA errors; see Figure 9 in Section 5.4. A pragmatic solution is to use a low-frequency cutoff in the estimation of ARFIMA models in the frequency domain; however, this incurs a substantial loss of information in the estimation of the long-memory parameter $d$. Section 5 thus uses the exact Gaussian likelihood for dynamic linear models with ARFIMA errors, which corresponds to the exact maximum likelihood approach in Doornik and Ooms, (2004). The resulting computation is much slower than a frequency domain approach and stresses the importance of our contribution, namely: a computationally fast dynamic linear model that accounts for long-range dependence in the error process.
@@ 5.1. Overview and objectives
This section analyses electricity demand as the response variable conditional on relevant exogenous variables for two real-world datasets. The two datasets contain data for New England, United States of America (Example 1), and Victoria, Australia (Example 2), presented in Sections 5.4 and 5.5, respectively. Forecasting for Example 2 is considerably more challenging than for Example 1 and we show that accounting for long-range dependence is essential to obtain good forecast accuracy for moderately long horizons. In contrast, long-range dependency does not improve the forecasts in Example 1; however, we show it has a significant impact on the posterior distribution of the regression parameters $\boldsymbol{\beta}$.
@@ 5.2. Model settings
We consider three types of error processes: ARTFIMA, ARFIMA and ARMA. To choose the order of $p$ and $q$, a search over the model space is considered up to a maximum of $p=2,q=1$. For various orders of $p$ and $q$, we select the model with the smallest deviance information criterion (DIC) value (Spiegelhalter et al.,, 2002). For a given model $\mathcal{M}$, the deviance is defined as
$D(\boldsymbol{\theta})=-2\log p(\boldsymbol{y}|\boldsymbol{\theta},\mathcal{M})+2\log h(\boldsymbol{y}),$ (5.1)
where $h(\boldsymbol{y})$ is a fully specified standardising term which is a function of only the data, which we set to be $h(\boldsymbol{y})=1$ for all models. The DIC consists of two components,
$DIC=\overline{D}(\boldsymbol{\theta})+p_{D},$ (5.2)
where the first term is the posterior expectation of the deviance in (5.1),
$\overline{D(\boldsymbol{\theta})}=\text{E}\left[D(\boldsymbol{\theta})\right]=\text{E}\left[-2\log p(\boldsymbol{y}|\boldsymbol{\theta},\mathcal{M})\right].$
The first term in (5.2) assesses how well the model fits the data, with smaller values being preferred. The second component, $p_{D}$, is the effective number of parameters of the model, which penalises the complexity of the model,
$p_{D}=\overline{D(\boldsymbol{\theta})}-D({\boldsymbol{\theta}}^{\star})=\text{E}\left[-2\log p(\boldsymbol{y}|\boldsymbol{\theta},\mathcal{M})\right]+2\log p(\boldsymbol{y}|{\boldsymbol{\theta}}^{\star},\mathcal{M}),$
where ${\boldsymbol{\theta}}^{\star}$ is the posterior mean. Thus, the DIC is a trade-off between the complexity and the adequacy of a model (Chan and Grant,, 2016).
Once the best $p$ and $q$ are chosen for the ARMA error process, we compare the ARMA model to the ARTFIMA and ARFIMA models with the same $p$ and $q$. Likewise, we find the best $p$ and $q$ for the ARTFIMA error process and compare it with the ARMA and ARFIMA models with the same $p$ and $q$. However, this is not done for ARFIMA since performing MCMC over the whole model space under the Gaussian likelihood is impractical with the large datasets in the applications.
@@
It is well known (e.g. Wheeler and Ionides,, 2024) that the log-likelihood function for ARMA models may exhibit multimodality, for example when root or near root cancellation is present, and this is also true for ARTFIMA and ARFIMA. Moreover, the parameters $d$ and $\lambda$ introduce additional complexity to the geometry of the likelihood function. To avoid finding a poor local mode, we use global optimisation via Basin-hopping (Li and Scheraga,, 1987) to find the MAP estimate, which serves as an initial starting value for MCMC.
5.3. Settings for forecast evaluation
For a given $p$ and $q$ determined as described in Section 5.2, we perform time series leave-future-out-cross-validation to evaluate the forecasting performance. We follow Hyndman and Athanasopoulos, (2021, Ch. 7, Ch. 10) and assume all exogenous variables $\mathbf{X}_{t}$ are known when forecasting. We use a sliding/rolling window approach with a fixed training size of $T$ with $k=100$ out-of-sample observations for testing. To compare models, we use three different metrics: log-predictive density score (LPDS), root mean square error (RMSE) and continuous rank probability score (CRPS). We compute each metric for different forecast horizons, $h=1,\dots,15$ and note that the $h$-step ahead case results in $k-h+1$ out-of-sample observations ($k=100$). Each time the window is rolled forward (to predict the next out-of-sample observation), we re-estimate the model and forecast $h$-steps-ahead. However, for ARFIMA, the model is only estimated once on the initial training set; this is because re-estimating the model is very costly due to the evaluation of the Gaussian likelihood. This does not have a large impact since the sliding window approach only impacts the posterior negligibly due to the large amount of data. We note that approximate leave-future-out cross-validation techniques have been proposed (Bürkner et al.,, 2020) for computing the LPDS that do not require reestimating the posterior. The other metrics used in this paper, however, require the posterior samples and the cost of computing the LPDS is negligible given these samples. Hence we do not implement Bürkner et al., (2020).
The $h$-step-ahead log-predictive density score (LPDS) is
$\text{LPDS}^{(h)}=\frac{1}{k-h+1}\sum^{k-h}_{i=0}\log p(y_{T+h+i}|y_{1+i:T+i}),$
where the $y_{T+h+i}$ for $i=0,\dots,k-h$ are the out-of-sample observations for which we evaluate the log-predictive density,